Two cyclists start 60km apart. If they cycle towards each other, they meet in an hour and if they cycle in the same direction they meet in five hours. Find their average speeds. (Hint: Draw a sketch)

Answer by srinivas.g(540)

Distance = 60 km
let x and y be the average speeds of two cycles
if two cycles travel in opposite direction, then relative speed = x+y
if two cycles travel in same direction , relative speed = x-y
but if two cycles met in an hour in opposite direction. Hence they covered a distance of 60 km in 1 hour. hence their relative speed is 60 km per hour
i.e x+y = 60 kmph
if they travel in same direction , they met in five hours
hence relative speed (x-y) = 60 km /5 hrs
= 12 Kmph
x-y = 12 kmph
solving both equations
x+y = 60
x-y = 12
adding both equations
2x =72
x= 36 kmph
y = 36-12
=24 kmph
so speed of both cycles will be 36 kmph and 24 kmph